Find the equation of the plane which passes through the line of intersection of the planes x+y-2z=1 and 2x+y-4z=3 and which is perpendicular to the plane x+y+z=1.
1
Expert's answer
2019-07-16T10:48:45-0400

Solution:

Equation of the plane passing through the line of intersection of the plane


A1x + B1y + C1z = d1 and -------------------- (1)


A2x + B2y + C2z = d2 -------------------- (2)


(A1x + B1y + C1z - d1 ) + "\lambda" ( (A2x + B2y + C2z - d2 ) ------------------ (G)


Given equation of the planes


x+y-2z=1 ----------------------- (3)


and 2x+y-4z=3 ----------------------- (4)


Comparing equation (1) and (3)


A1 = 1 , B1 = 1 , C1 = -2 , d1 = 1


Comparing equation (2) and (4)


A2 = 2, B2= 1 , C2 = -4 , d2 = 3


Putting the values in equation (G)


( x + y -2z -1 ) + λ(2x + y -4z -3 )


x + y -2z -1 + 2λx + λy -4λz -3λ = 0


(1+2λ)x + (1+λ)y + (-2 -4λ) z + (-1-3λ) = 0 ---------------- (5)


Also, the plane is perpendicular to the plane x + y + z = 1 ------------- (6)


So the normal vector "\overrightarrow{N}" to be the plane is perpendicular to the normal vector of x + y + z =1


"\overrightarrow{N}" = (1+2λ)"\widehat{i}" + (1+λ)"\widehat{j}" + (-2 -4λ) "\widehat{j}" --------------------- (7)


and "\overrightarrow{n}" = 1"\widehat{i}" + 1"\widehat{j}" + 1"\widehat{k}" ----------------------(8)


We know that two lines with direction ratio a1, b1,c1 and a2,b2,c2 are perpendicular if


a1a2 + b1b2 + c1c2 = 0 ----------------------(9)


From equation (7) and (8) we have


a1= (1+2λ)

b1 = (1+λ)

c1 = (-2 -4λ)

a2 = 1

b2 = 1

c2 = 1


Putting the values in equation (9)

(1+2λ) (1) + (1+λ) (1) + (-2 -4λ)(1) = 0

1+2λ + 1+λ -2 -4λ = 0

3λ - 4λ = 0

So we get λ = 0


Put the value of λ in equation (5)

(1+2λ)x + (1+λ)y + (-2 -4λ) z + (-1-3λ) = 0


(1+2*0)x + (1+0)y + (-2 -4*0) z + (-1-3*0) = 0


(1)x + (1)y +(-2)z -1 = 0


x + y-2z -1 =0 (Answer)